OpenStudy (anonymous):

g(x) = 4x(x^2 − 4x − 3) (a) Find all the real zeros of the polynomial function.

OpenStudy (alexwee123):

can't you factor out (x^2 − 4x − 3)?

sam (.sam.):

Set it =0 and solve

OpenStudy (anonymous):

I"m having a difficult time with it for some reason. it's not coming to my head.

OpenStudy (mertsj):

Well, the trinomial will not factor so you will have to use the quadratic formula to find the zeros of that factor.

sam (.sam.):

\[4 x \left(x^2-4 x-3\right)=0\]

OpenStudy (mertsj):

So you have 4x=0 or x^2-4x-3=0 x=0 and solve the quadratic with the formula.

sam (.sam.):

\[x=0\text{ or }4 x^2-16 x-12=0\]

OpenStudy (anonymous):

OH RIGHT. i totally forgot about the quad formula, lol.

OpenStudy (mertsj):